. Consider a batch experiment during which the aerobic consumption of glucose is tracked in terms of the associated oxygen consumption. Given an initial glucose concentration of 100 mg COD/L and a final cumulative oxygen consumption of 75 mg O2/L

a. Calculate the biomass yield coefficient, expressed as mg COD biomass produced per mg COD glucose consumed

b. Determine the amount of biomass produced towards the end of the assay neglecting biomass decay
Answer #1

Here we first need to find the COD/mole of Glucose and Biomass

Complete oxidation of glucose by oxygen can be given by following equation-

C6H12O6 + 6O2 = 6CO2 + 6H2O

From above balanced equation we can calculate COD for glucose (mol wt 180) as

6moles*32g/mole Oxygen/1mole*180g/mole = 1.06 O2 / Glucose.

Now we know that the empirical formula for biomass is C5H7O2N

Complete oxidation of biomass by oxygen can be given by following equation-

C5H7O2N + 5O2 = 5CO2 + 2H2O +NH3

From above balanced equation we can calculate COD for biomass (mol wt 113) as

5moles*32g/mole Oxygen/1mole*113g/mole = 1.416 O2 / Biomass

a) Now, balanced equation for aerobic conversion of glucose into biomass is-

C6H12O6 + O2 + NH3 = C5H7O2N + CO2 + 4H2O…..(1)

From above balanced equation we know

Biomass yield = 1 mole of biomass/ 1 mole of glucose.

Now we convert the above yields in COD of biomass and glucose

Therefore biomass yield coefficient = 1.416 mg COD of biomass/ 1.06 mg COD of glucose

                                                                 = 1.33 mg COD of biomass/mg COD of glucose

b) From equation (1) we know 1 mole of biomass is produced by consumption of 1 mole of oxygen. We can also say that 113 g of biomass is produced by consumption of 32 g of oxygen.

Now, it is given in the question that 75 mg O2/L is consumed during the assay. Therefore total biomass produced towards the end of the assay can be calculated as-

113*75/32 = 264.84 mg biomass/L


Thanks!